# New versions of Frobenius and integral closure of ideals

### v2

@@ 3.4. Comparing the new closures
First note that when $\mathfrak{a}=R,$ the above result recovers the usual Briançon-Skoda theorem for tight closure. Further, a similar Briançon-Skoda theorem holds for the Vraciu $\mathfrak{a}^{t}$-integral and tight closures, but since ${}^{\mathfrak{a}^{t}}\overline{I}=\overline{I}$, the result is true simply by the usual Briançon-Skoda; ${}^{\mathfrak{a}^{t}}\overline{I^{m+n}}=\overline{I^{m+n}}\subset(I^{m+1})^{*}\subset$ (${}^{\mathfrak{a}^{t}}(I^{m+1})^{*}.$
To conclude this section, we say a few words about these adjusted closures in the graded setting and prove a graded version of the Briançon-Skoda Theorem.
Setting 3.21.
@@
Given 3.21, let $I=(f_{1},\ldots,f_{d})$ be a homogeneous ideal, and $x\in R_{\geq N}$ with $N=\beta k-\beta^{\prime}l+\max(\deg(f_{i})).$
@@
    Then, $x\in\overline{I}^{\mathfrak{a}}$ if and only if $x\in$ ${}^{\mathfrak{a}}\overline{I}$ = $\overline{I}$, and
@@
    $x\in I^{F\mathfrak{a}}$ if and only if $x\in$ ${}^{\mathfrak{a}}I^{F}.$
@@
The proof of the second claim is a quick adaptation of the proof of [Vra08, Proposition 2.9], by simply removing the multiplier $c$ from the equations involved. We now turn to the first claim.
Assume that $x\in\overline{I}^{\mathfrak{a}}$ and $\deg(x)\geq N.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{n}\subset I^{n}$ for all $n$ large. Let $h\in\mathfrak{a}^{n}$ be a homogeneous element. Then $\deg(h)\geq\beta^{\prime}ln.$ Write $cx^{n}h=\sum_{|\mathbf{m}|=n}b_{\mathbf{m}}f_{1}^{m_{1}}\cdots f_{d}^{m_{d}}$ where $b_{i}\in R$ are homogeneous elements, so that
$\deg(b_{\mathbf{m}})\geq\deg(c)+n\deg(x)+\deg(h)-\max(\deg(f_{i}))n\geq\beta kn,$
for each $\mathbf{m}.$ This implies that $b_{\mathbf{m}}\in R_{\geq\beta kn}\subset R_{+}^{kn}\subset\mathfrak{a}^{n},$ hence the result. ∎
@@
Given 3.21, suppose $\operatorname{char}\mathbb{F}=p>0$ and $R$ is locally equidimensional. Let $I=(f_{1},\ldots,f_{d})$ be such that $f_{1},\ldots,f_{d}$ are homogeneous system of parameters, and let $N=\sum_{i=1}^{d}\deg f_{i}-\beta^{\prime}l.$
Suppose $R$ has a parameter test element [^2] Recall an element $c\in R^{\circ}$ is a parameter test element if, for all $e\geq 0$ and all ideals $\mathfrak{q}$ generated by a system of parameters of $R$, $c(\mathfrak{q}^{*})^{\left[p^{e}\right]}\subset\mathfrak{q}^{\left[p^{e}\right]}$., e.g. $\mathbb{F}$ is $F$-finite. Let $x\in R_{\geq N}.$ Then $x\in I^{*\mathfrak{a}}.$ In other words, $R_{\geq N}+I^{F\mathfrak{a}}\subset I^{*\mathfrak{a}}.$
@@
We first prove the case in which $\deg f_{i}=\deg f_{1}=\alpha$ for all $1\leq i\leq d.$ Pick $v\in\mathfrak{a}$ a homogeneous element. As $\deg v\geq\beta^{\prime}l,$ we get $\deg(xv)\geq d\alpha.$ Since $R$ is module-finite over $S=\mathbb{F}[f_{1},\ldots,f_{d}],$ $xv$ satisfies an integral equation $(xv)^{m}+b_{1}(xv)^{m-1}+\cdots+b_{m}=0$ where $b_{i}\in S$ for all $i.$ We may assume that $b_{i}$ are homogeneous elements. So $\deg b_{i}=i\cdot\deg(xv)$ implying that
$b_{i}\in R_{i\cdot\deg(xv)}\cap S\subset(f_{1},\ldots,f_{d})^{i\cdot\deg(xv)/\alpha}\subset(f_{1},\ldots,f_{d})^{id}.$
@@
This implies that $xv\in\overline{I^{d}}.$ By definition, there exists $c_{0}\in R^{\circ}$ and $n_{0}$ such that $c_{0}x^{n}v^{n}\in I^{nd},$ for all $n\geq n_{0}.$ Note that $I^{n_{0}}\cap R^{\circ}\neq\varnothing.$ Choose $c_{1}\in I^{n_{0}}\cap R^{\circ}$ and set $c=c_{0}c_{1}.$ Then for all $n\geq 0,$
(3.1) $cx^{n}v^{n}\in I^{nd}.$

### v3

@@ 4.1. Briançon-Skoda theorems
[…]
4.2. Comparisons with sharp $F$-closure
In [Sch08], Schwede defined the $\mathfrak{a}^{t}$-sharp Frobenius closure of $I$, denoted $I^{F^{\sharp}\mathfrak{a}^{t}}$, to be the set of all $x\in R$ such that for all $e\gg 0$, $x^{p^{e}}\mathfrak{a}^{\lceil t(p^{e}-1)\rceil}\subset I^{[p^{e}]}$. It is clear that the $\mathfrak{a}^{t}$-sharp Frobenius closure is only a preclosure; $\bullet^{F^{\sharp}\mathfrak{a}^{t}}$ is extensive and order-preserving by a similar proof to Proposition 3.15, but taking the example $R=\mathbb{F}_{p}[x]$ and $I=\mathfrak{a}=(x^{2})$, one can easily show $(I^{F^{\sharp}\mathfrak{a}})^{F^{\sharp}\mathfrak{a}}$ is strictly larger than $I^{F^{\sharp}\mathfrak{a}}$.
Schwede’s definition differs from our Hara-Yoshida $\mathfrak{a}^{t}$-Frobenius closure due to the smaller exponent on $\mathfrak{a}$, as the following example shows. Note that by definition we have that $I^{F^{\sharp}\mathfrak{a}^{t}}\subset I^{F\mathfrak{a}^{t}}$.
Example 4.7.
Let $R=\mathbb{F}_{p}[x,y]$, $I=(x,y)$, $\mathfrak{a}=(x,y)$, and $t=2$. We claim that $1\in I^{F\mathfrak{a}^{t}}\setminus I^{F^{\sharp}\mathfrak{a}^{t}}$. First, note that for any $e\gg 0$, $1\mathfrak{a}^{\lceil tp^{e}\rceil}=(x,y)^{2p^{e}}\subset(x^{p^{e}},y^{p^{e}})=I^{[p^{e}]}$, so $1\in I^{F\mathfrak{a}^{t}}$. On the other hand, $1\mathfrak{a}^{\lceil t(p^{e}-1)\rceil}=(x,y)^{2(p^{e}-1)}$ contains $x^{p^{e}-1}y^{p^{e}-1}\notin I^{[p^{e}]}$. Therefore $1\notin I^{F^{\sharp}\mathfrak{a}^{t}}$.
4.7 shows that a small adjustment in the exponent of $\mathfrak{a}$ in the definition of the adjusted Frobenius closure can give a different operation on ideals. The following example shows that adjusting the exponent of $\mathfrak{a}$ in the definition of the Vraciu Frobenius closure also provides a different operation on ideals.
Example 4.8.
Let $R=\mathbb{F}_{p}[x,y]$, $I=(x^{2},y^{2})$, $\mathfrak{a}=(x,y)$, and $t=2$. First, note that $^{\mathfrak{a}^{t}}I^{F}$ contains the set $\left\{x\in R\mid\forall e\gg 0,x^{p^{e}}\mathfrak{a}^{\lceil t(p^{e}-1)\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil t(p^{e}-1)\rceil}\right\}$.
Now, for any $e\geq 0$, the ideal $(xy)^{p^{e}}\mathfrak{a}^{2p^{e}}$ is generated by monomials of the form $x^{p^{e}+i}y^{3p^{e}-i}$ for $0\leq i\leq 2p^{e}$. If $i<p^{e}$, then $x^{p^{e}+i}y^{3p^{e}-i}=y^{2p^{e}}x^{p^{e}+i}y^{p^{e}-i}\subset I^{[p^{e}]}\mathfrak{a}^{2p^{e}}$, and if $i\geq p^{e}$ then $x^{p^{e}+i}y^{3p^{e}-i}=x^{2p^{e}}x^{i-p^{e}}y^{3p^{e}-i}\subset I^{[p^{e}]}\mathfrak{a}^{2p^{e}}$. Thus $(xy)^{p^{e}}\mathfrak{a}^{2p^{e}}\subset I^{[p^{e}]}\mathfrak{a}^{2p^{e}}$, and so $xy\in{}^{\mathfrak{a}^{t}}I^{F}$. On the other hand, it is not the case that $(xy)^{p^{e}}\mathfrak{a}^{2(p^{e}-1)}\subset I^{[p^{e}]}\mathfrak{a}^{2(p^{e}-1)}$. The former set contains the monomial $(xy)^{p^{e}}x^{p^{e}-1}y^{p^{e}-1}=x^{2p^{e}-1}y^{2p^{e}-1}$, which is not even contained in $I^{[p^{e}]}$. Therefore,
$\left\{x\in R\mid\forall e\gg 0,x^{p^{e}}\mathfrak{a}^{\lceil t(p^{e}-1)\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil t(p^{e}-1)\rceil}\right\}\subsetneq{}^{\mathfrak{a}^{t}}I^{F}.$
By contrast, the other adjusted closure operations, by virtue of including the multiplier $c\in R^{\circ}$ in their definitions, are stable under modification of the exponent on $\mathfrak{a}$.
Theorem 4.9.
Let $u\in\mathbb{R}$. If $R$ is a ring, $I,\mathfrak{a}$ are ideals of positive height, and $t\in\mathbb{R}_{>0}$, then
  1. (a)
    $\overline{I}^{\mathfrak{a}^{t}}=\left\{x\in R\mid\text{ for some }c\in R^{\circ}\text{ and all }n\gg 0,cx^{n}\mathfrak{a}^{\lceil tn+u\rceil}\subset I^{n}\right\}$.
  2. (b)
    $I^{*\mathfrak{a}^{t}}=\left\{x\in R\mid\text{ for some }c\in R^{\circ}\text{ and all }e\gg 0,cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}+u\rceil}\subset I^{[p^{e}]}\right\}$.
  3. (c)
    $^{\mathfrak{a}^{t}}I^{*}=\left\{x\in R\mid\text{ for some }c\in R^{\circ}\text{ and all }e\gg 0,cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}+u\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}+u\rceil}\right\}$.
Proof.
(a) Let $K_{u}=\left\{x\in R\mid\text{ for some }c\in R^{\circ}\text{ and all }n\gg 0,cx^{n}\mathfrak{a}^{\lceil tn+u\rceil}\subset I^{n}\right\}$. We wish to show that $K_{u}=\overline{I}^{\mathfrak{a}^{t}}$.
Suppose $u>0$, so that $\mathfrak{a}^{\lceil tn+u\rceil}\subset\mathfrak{a}^{\lceil tn\rceil}$ and hence $\overline{I}^{\mathfrak{a}^{t}}\subset K_{u}$. Now suppose that $x\in K_{u}$, let $c^{\prime}\in R^{\circ}$ such that $c^{\prime}x^{n}\mathfrak{a}^{\lceil tn+u\rceil}\subset I^{n}$ for $n\gg 0$, and let $c^{\prime\prime}\in\mathfrak{a}^{\lceil u\rceil}\cap R^{\circ}$. Now, for $n\gg 0$,
$c^{\prime}c^{\prime\prime}x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset c^{\prime}x^{n}\mathfrak{a}^{\lceil tn\rceil+\lceil u\rceil}\subset c^{\prime}x^{n}\mathfrak{a}^{\lceil tn+u\rceil}\subset I^{n},$
and so $x\in\overline{I}^{\mathfrak{a}^{t}}$.
On the other hand, if $u<0$, then a priori $K_{u}\subset\overline{I}^{\mathfrak{a}^{t}}$, so let $x\in\overline{I}^{\mathfrak{a}^{t}}$, with a choice of $c^{\prime}\in R^{\circ}$ such that $c^{\prime}x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n\gg 0$. Choose $c^{\prime\prime}\in\mathfrak{a}^{\lceil-u\rceil}\cap R^{\circ}$, and note that for $n\gg 0$,
$c^{\prime}c^{\prime\prime}x^{n}\mathfrak{a}^{\lceil tn+u\rceil}\subset c^{\prime}x^{n}\mathfrak{a}^{\lceil tn+u\rceil+\lceil-u\rceil}\subset c^{\prime}x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n},$
hence $x\in K_{u}$.
(b) Replacing $I^{n}$ with $I^{[p^{e}]}$ in the previous two paragraphs shows that the same result holds for Hara-Yoshida tight closure.
(c) Let $L_{u}=\left\{x\in R\mid\text{ for some }c\in R^{\circ}\text{ and all }e\gg 0,cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}+u\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}+u\rceil}\right\}$. We wish to show that $L_{u}={}^{\mathfrak{a}^{t}}I^{*}$.
First suppose $u>0$. Since for all $e\geq 0$, $(I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil}:\mathfrak{a}^{\lceil tp^{e}\rceil})\subset(I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}+u\rceil}:\mathfrak{a}^{\lceil tp^{e}+u\rceil})$, we have that $^{\mathfrak{a}^{t}}I^{*}\subset L_{u}$. Suppose that $x\in L_{u}$ and choose $c^{\prime}\in R^{\circ}$ such that $c^{\prime}x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}+u\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}+u\rceil}$ for all $e\gg 0$. Choose $c^{\prime\prime}\in\mathfrak{a}^{\lceil u\rceil}\cap R^{\circ}$ and note that
$c^{\prime}c^{\prime\prime}x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset c^{\prime}x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}+u\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}+u\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil},$
and so $x\in{}^{\mathfrak{a}^{t}}I^{*}$.
If $u<0$, then a priori $L_{u}\subset{}^{\mathfrak{a}^{t}}I^{*}$. Let $x\in{}^{\mathfrak{a}^{t}}I^{*}$, let $c^{\prime}\in R^{\circ}$ such that for $e\gg 0$, $c^{\prime}x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil}$. Choose $c^{\prime\prime}\in\mathfrak{a}^{\lceil-u\rceil}\cap R^{\circ}$, and note that
$c^{\prime}c^{\prime\prime}x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}+u\rceil}\subset c^{\prime}x^{n}\mathfrak{a}^{\lceil tp^{e}+u\rceil+\lceil-u\rceil}\subset c^{\prime}x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}+u\rceil},$
hence $x\in L_{u}$. ∎
We also note that we have a similar Briançon-Skoda type containment between the $\mathfrak{a}^{t}$-integral closure and sharp $F$-closure.
Proposition 4.10.
Let $R$ be a ring and $\mathfrak{a},I\subset R$ be ideals of positive height. Let $I$ be generated by $d$ elements and let $t\in\mathbb{R}_{>0}$. Then $\overline{I^{d+m}}^{\mathfrak{a}^{t}}\subset(I^{m})^{F^{\sharp}\mathfrak{a}^{t}}$ for all $m\geq 1$.
Proof.
Let $x\in\overline{I^{d+m}}^{\mathfrak{a}^{t}}$. As in the proof of Theorem 4.5(b), we may first assume that $R$ is reduced and that there exists $n^{\prime}$ such that for all $n\gg 0$, $x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n-n^{\prime}}$. Therefore, for $e\gg 0$,
$x^{p^{e}}\mathfrak{a}^{\lceil t(p^{e}-1)\rceil}\subset x^{p^{e}-1}\mathfrak{a}^{\lceil t(p^{e}-1)\rceil}\subset I^{(d+m)(p^{e}-1)-n^{\prime}}\subset(I^{[p^{e}]})^{\lceil m-(d+m+n^{\prime})/p^{e}\rceil}=(I^{[p^{e}]})^{m}=(I^{m})^{[p^{e}]}.$
Therefore $x\in(I^{m})^{F^{\sharp}\mathfrak{a}^{t}}$. Generally, if $R$ is not reduced, then from the previous case, we know that for all $m\geq 1,$
$\overline{I^{d+m}}^{\mathfrak{a}^{t}}=\overline{I^{d+m}}^{\mathfrak{a}^{t}}+\sqrt{(0)}\subset(I^{m})^{F^{\sharp}\mathfrak{a}^{t}}+\sqrt{(0)}=(I^{m})^{F^{\sharp}\mathfrak{a}^{t}},$
where the last equality follows from the observation that $\sqrt{(0)}=(0)^{F}\subset(I^{m})^{F^{\sharp}\mathfrak{a}^{t}}.$
4.3. Adjusted closures in graded rings
To conclude this section, we say a few words about these adjusted closure operations in a graded setting and prove a graded version of the Briançon-Skoda Theorem. First, we note that in $\mathbb{Z}^{N}$-graded rings, the adjusted closure operations applied to homogeneous ideals return homogeneous ideals. If $R$ is $\mathbb{Z}^{N}$-graded, for an $x=\sum_{\mathbf{j}\in\mathbb{Z}^{N}}x_{\mathbf{j}}$ in $R$, write $\operatorname{Supp}x=\{\mathbf{j}\in\mathbb{Z}^{N}\mid x_{\mathbf{j}}\neq 0\}$. Furthermore, we will endow $\mathbb{Z}^{N}$ with a total ordering $\leq$ which respects addition, e.g. the lexicographic order. The key idea of the proof below is to use that the equations defining membership in the adjusted closures tend to isolate elements of maximal graded degree with respect to $\leq$.
Proposition 4.11.
Suppose $R=\oplus_{\mathbf{m}\in\mathbb{Z}^{N}}R_{\mathbf{m}}$ is a $\mathbb{Z}^{N}$-graded ring, $I$ and $\mathfrak{a}$ are homogeneous ideals of positive height in $R$, and $t\in\mathbb{R}_{>0}$. Then, the $\mathfrak{a}^{t}$-integral closure $\overline{I}^{\mathfrak{a}^{t}}$ is also a homogeneous ideal. If $R$ is of prime characteristic $p>0$, then each of the adjusted closures $I^{F\mathfrak{a}^{t}}$, $I^{*\mathfrak{a}^{t}}$, $^{F}I^{\mathfrak{a}^{t}}$, and $^{*}I^{\mathfrak{a}^{t}}$ are homogeneous ideals as well.
Proof.
Define $I_{n}=I^{n}:\mathfrak{a}^{\lceil tn\rceil}$, $J_{e}=I^{\left[p^{e}\right]}:\mathfrak{a}^{\lceil tp^{e}\rceil}$, and $J_{e}^{\prime}=I^{\left[p^{e}\right]}\mathfrak{a}^{\lceil tp^{e}\rceil}:\mathfrak{a}^{\lceil tp^{e}\rceil}$. Since products and colons of homogeneous ideals are homogeneous, we have that $I_{n}$, $J_{e}$, and $J_{e}^{\prime}$ are homogeneous ideals for all $n,e\in\mathbb{N}$.
We first handle the adjusted integral closure. Let $x\in R$ be an element of $\overline{I}^{\mathfrak{a}^{t}}$ so that there is a $c\in R^{\circ}$ such that $cx^{n}\in I_{n}$ for all $n$. By 3.8, we can assume without loss of generality that $R$ is a domain. Our goal is that $x_{\mathbf{j}}\in\overline{I}^{\mathfrak{a}^{t}}$ for all $\mathbf{j}\in\operatorname{Supp}x$. We will induce on the size of $\operatorname{Supp}x$. If $x$ itself is homogeneous, then there’s nothing to show. Now, suppose that if $|\operatorname{Supp}y|=m$ and $cy^{n}\in I_{n}$ for all $n$, then each homogeneous component of $y$ is in $\overline{I}^{\mathfrak{a}^{t}}$, and now suppose $x$ has $|\operatorname{Supp}x|=m+1$.
Let $\mathbf{j}_{0}$ be the maximum element of $\operatorname{Supp}x$ under $\leq$ and let $\mathbf{i}_{0}$ be the maximal element of $\operatorname{Supp}c$ under $\leq$. Then, an arbitrary homogeneous term of the multinomial expansion (suppressing the multinomial coefficient) of $c(\sum_{\mathbf{j}\in\mathbb{Z}^{N}}x_{\mathbf{j}})^{n}$, has the form $c_{\mathbf{i}}x_{\mathbf{j}_{0}}^{e_{0}}x_{\mathbf{j}_{1}}^{e_{1}}\cdots x_{\mathbf{j}_{m}}^{e_{m}}$ where $\sum_{k}e_{k}=n$. This element has degree
$\mathbf{i}+\sum_{k}e_{k}\mathbf{j}_{k}\leq\mathbf{i}_{0}+\sum_{k}e_{k}\mathbf{j}_{0}=\mathbf{i}_{0}+n\mathbf{j}_{0}$
with equality if and only if $\mathbf{i}=\mathbf{i}_{0}$ and $e_{k}=0$ for $k>0$. Hence, $c_{\mathbf{i}_{0}}x_{\mathbf{j}_{0}}^{n}$ is the only homogeneous term of $cx^{n}$ of graded degree $\mathbf{i}_{0}+n\mathbf{j}_{0}$. Thus, as $I_{n}$ is homogeneous, $c_{\mathbf{i}_{0}}x_{\mathbf{j}_{0}}^{n}\in I_{n}$ for all $n$. Hence $x_{\mathbf{j}_{0}}\in\overline{I}^{\mathfrak{a}^{t}}$, and so $x-x_{\mathbf{j}_{0}}\in\overline{I}^{\mathfrak{a}^{t}}$ as well. By the induction hypothesis, each homogeneous component of $x-x_{\mathbf{j}_{0}}$ is in $\overline{I}^{\mathfrak{a}^{t}}$, so each homogeneous component of $x$ is in $\overline{I}^{\mathfrak{a}^{t}}$, concluding the proof.
Now we suppose that $R$ has prime characteristic $p>0$, and we discontinue the supposition that $R$ is a domain since the Frobenius closures cannot be computed by passing to quotient domains.
As before, fix $x=\sum_{\mathbf{j}\in\mathbb{Z}^{N}}x_{\mathbf{j}}$ and $c=\sum_{\mathbf{i}\in\mathbb{Z}^{N}}c_{\mathbf{i}}\in R^{\circ}$. Write $\mathbf{j}_{0}$ for the maximal element of $\operatorname{Supp}x$ under $\leq$. Then, for $\mathbf{j}\neq\mathbf{j}_{0}$, we have that the support of $cx_{\mathbf{j}}^{p^{e}}$ has maximum $\mathbf{i}_{0}+p^{e}\mathbf{j}$, which for $e\gg 0$ is less than $\mathbf{i}+p^{e}\mathbf{j}_{0}$ for all $\mathbf{i}\in\operatorname{Supp}c$. Hence, if $cx^{p^{e}}\in J_{e}$, then $cx_{\mathbf{j}_{0}}^{p^{e}}\in J_{e}$ as $J_{e}$ is homogeneous.
If $c=1$, this argument demonstrates that $x\in I^{F\mathfrak{a}^{t}}$ implies $x_{\mathbf{j}_{0}}\in I^{F\mathfrak{a}^{t}}$. If $c$ is simply an arbitrary element of $R^{\circ}$, this argument demonstrates that $x\in I^{*\mathfrak{a}^{t}}$ implies $x_{\mathbf{j}_{0}}\in I^{*\mathfrak{a}^{t}}$. Similar reasoning for membership in $J_{e}^{\prime}$ holds, showing if $x\in{}^{\mathfrak{a}^{t}}I^{*}$ (respectively $x\in{}^{\mathfrak{a}^{t}}I^{F}$) then $x_{\mathbf{j}_{0}}\in{}^{\mathfrak{a}^{t}}I^{*}$ (respectively $x_{\mathbf{j}_{0}}\in{}^{\mathfrak{a}^{t}}I^{F}$). Then, applying an induction argument on the size of $\operatorname{Supp}x$ as before, we see that $I^{*\mathfrak{a}^{t}}$, $I^{F\mathfrak{a}^{t}}$, $^{\mathfrak{a}^{t}}I^{*}$, and $^{\mathfrak{a}^{t}}I^{F}$ are homogeneous. ∎
We now switch to the more common setting of an $\mathbb{N}$-graded ring.
Setting 4.12.
@@ 4.3. Adjusted closures in graded rings
Given 4.12, let $I=(f_{1},\ldots,f_{d})$ be a homogeneous ideal, $t\in\mathbb{R}_{>0}$, and $x\in R_{\geq N}$ with $N=(\beta k-\beta^{\prime}l)t+\max(\deg(f_{i})).$
@@
    Then, $x\in\overline{I}^{\mathfrak{a}^{t}}$ if and only if $x\in$ $^{\mathfrak{a}^{t}}\overline{I}=\overline{I}$, and
@@
    If $t\in\mathbb{Z}[1/p]$ or $\deg(x)>N$, then $x\in I^{F\mathfrak{a}^{t}}$ if and only if $x\in$ $^{\mathfrak{a}^{t}}I^{F}.$
@@
(a) Assume that $x\in\overline{I}^{\mathfrak{a}^{t}}$ and $\deg(x)\geq N.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n$ large. Without loss of generality we may assume that $\deg(c)\geq\beta k-\beta^{\prime}l$. Fix $n$ and let $h\in\mathfrak{a}^{\lceil tn\rceil}$ be a homogeneous element. Then $\deg(h)\geq\beta^{\prime}l\lceil tn\rceil.$ Write $cx^{n}h=\sum_{|\mathbf{m}|=n}b_{\mathbf{m}}f_{1}^{m_{1}}\cdots f_{d}^{m_{d}}$ where $b_{i}\in R$ are homogeneous elements, so that
$\deg(b_{\mathbf{m}})$ $\geq\deg(c)+n\deg(x)+\deg(h)-n\max(\deg(f_{i}))$
$\geq\deg(c)+\beta ktn+\beta^{\prime}l(\lceil tn\rceil-tn)$
$=\deg(c)+\beta k\lceil tn\rceil-(\beta k-\beta^{\prime}l)(\lceil tn\rceil-tn)$
$\geq\beta k\lceil tn\rceil$
for each $\mathbf{m}.$ This implies that $b_{\mathbf{m}}\in R_{\geq\beta k\lceil tn\rceil}\subset R_{+}^{k\lceil tn\rceil}\subset\mathfrak{a}^{\lceil tn\rceil},$ hence the result.
(b) Suppose that $x\in I^{F\mathfrak{a}^{t}}$, so that for all $e\gg 0$, $x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}$. Let $\delta=\deg(x)-N$. Fix $e$ and let $h\in\mathfrak{a}^{\lceil tp^{e}\rceil}$ be a homogeneous element. Then $\deg(h)\geq\beta^{\prime}l\lceil tp^{e}\rceil.$ Write $x^{p^{e}}h=\sum_{i=1}^{d}b_{i}f_{i}^{p^{e}}$ where $b_{i}\in R$ are homogeneous elements, so that
$\deg(b_{i})$ $=p^{e}\deg(x)+\deg(h)-p^{e}\deg(f_{i})$
$\geq p^{e}\left((\beta k-\beta^{\prime}l)t+\max(\deg(f_{i}))+\delta\right)+\beta^{\prime}l\lceil tp^{e}\rceil-p^{e}\max(\deg(f_{i}))$
$=(\beta k-\beta^{\prime}l)tp^{e}+\delta p^{e}+\beta^{\prime}l\lceil tp^{e}\rceil$
$=\beta k\lceil tp^{e}\rceil-(\beta k-\beta^{\prime}l)(\lceil tp^{e}\rceil-tp^{e})+\delta p^{e}.$
If $t\in\mathbb{Z}[1/p]$ and $e$ is large enough that $tp^{e}\in\mathbb{Z}$, then
$\beta k\lceil tp^{e}\rceil-(\beta k-\beta^{\prime}l)(\lceil tp^{e}\rceil-tp^{e})+\delta p^{e}=\beta k\lceil tp^{e}\rceil+\delta p^{e}\geq\beta k\lceil tp^{e}\rceil.$
On the other hand, if $t\notin\mathbb{Z}[1/p]$, then $\delta>0$, and for all $e$ large enough that $\delta p^{e}\geq\beta k-\beta^{\prime}l$,
$\beta k\lceil tp^{e}\rceil-(\beta k-\beta^{\prime}l)(\lceil tp^{e}\rceil-tp^{e})+\delta p^{e}>\beta k\lceil tp^{e}\rceil-(\beta k-\beta^{\prime}l)+(\beta k-\beta^{\prime}l)=\beta k\lceil tp^{e}\rceil.$
In either case, $\deg(b_{i})\geq\beta k\lceil tp^{e}\rceil$, and so $b_{i}\in\mathfrak{a}^{\lceil tp^{e}\rceil}$, showing that $x\in{}^{\mathfrak{a}^{t}}I^{F}$. ∎
@@
Given 4.12, suppose $\operatorname{char}\mathbb{F}=p>0$ and $R$ is locally equidimensional. Let $I=(f_{1},\ldots,f_{d})$ be such that $f_{1},\ldots,f_{d}$ are homogeneous system of parameters, $t\in\mathbb{R}_{>0}$, and $N=\sum_{i=1}^{d}\deg f_{i}-t\beta^{\prime}l.$
Suppose $R$ has a parameter test element [^3] Recall an element $c\in R^{\circ}$ is a parameter test element if, for all $e\geq 0$ and all ideals $\mathfrak{q}$ generated by a system of parameters of $R$, $c(\mathfrak{q}^{*})^{\left[p^{e}\right]}\subset\mathfrak{q}^{\left[p^{e}\right]}$., e.g. $\mathbb{F}$ is $F$-finite. Let $x\in R_{\geq N}.$ Then $x\in I^{*\mathfrak{a}^{t}}.$ In other words, $R_{\geq N}+I^{F\mathfrak{a}^{t}}\subset I^{*\mathfrak{a}^{t}}.$
@@
By 3.13, we may assume that $R$ is a domain. We first prove the case in which $\deg f_{i}=\deg f_{1}=\alpha$ for all $1\leq i\leq d$, so that $N=d\alpha-t\beta^{\prime}\ell$. Pick $0\neq v\in\mathfrak{a}$ a homogeneous element. For any $r\in\mathbb{N}$, since $R$ is module-finite over $S=\mathbb{F}[f_{1},\ldots,f_{d}],$ the monomial $x^{r}v^{\lfloor rt\rfloor}$ satisfies an integral equation $(x^{r}v^{\lfloor rt\rfloor})^{m}+b_{1}(x^{r}v^{\lfloor rt\rfloor})^{m-1}+\cdots+b_{m}=0$ where $b_{i}\in S$ for all $i.$ We may assume that $b_{i}$ are homogeneous elements. So for any $b_{i}\neq 0$,
$\deg b_{i}=i\cdot\deg(x^{r}v^{\lfloor rt\rfloor})\geq i\cdot\left(rN+\lfloor rt\rfloor\beta^{\prime}\ell\right)=i\cdot\left(rd\alpha-rt\beta^{\prime}\ell+\lfloor rt\rfloor\beta^{\prime}\ell\right).$
We will fix $r$ such that $rt\beta^{\prime}\ell-\lfloor rt\rfloor\beta^{\prime}\ell<1$, so that the inequality above will imply that $\deg b_{i}\geq i\cdot rd\alpha$. If $t$ is rational, then take $r$ such that $rt\in\mathbb{Z}$, so that $\lfloor rt\rfloor=rt$. Otherwise, if $t$ is irrational, then we may take $r\in\mathbb{N}$ such that $rt-\lfloor rt\rfloor<1/(\beta^{\prime}\ell)$. Hence
$b_{i}\in R_{i\cdot\deg(x^{r}v^{\lfloor rt\rfloor})}\cap S\subset(f_{1},\ldots,f_{d})^{i\cdot\deg(x^{r}v^{\lfloor rt\rfloor})/\alpha}\subset(f_{1},\ldots,f_{d})^{i\cdot rd}.$
@@
This implies that $x^{r}v^{\lfloor rt\rfloor}\in\overline{I^{rd}}.$ Thus there exists $c\in R^{\circ}$ such that $c(x^{r}v^{\lfloor rt\rfloor})^{n}\in I^{nrd}$ for all $n\geq 0.$